In tube #1, the original concentration of penicillin is 1000 units per ml. In a 1/5 dilution series, the first tube to show growth is tube #7. What is the m nmal hibitory concentration of penicillin in this assay? Given a plate with 58 colonies, determine the original concentration of bacteria in tube #1 for each of out 1.0 ml fontube #4 of a 1/10 dilution series, c) if one has plated out 0.1 ml from tube #5 of a l/10 dilution series; d) if one has plated out 025 ml from tube #5 of a l/10 dilution series. 5. 6 a) if one has plated out1
Microbiology
Answer #1

5. The tube 1 has 1000 units/ml.

1: 5 dilution of Tube 2= 1 ml of 1000 units/ml + 4 ml D/W

Concentration of tube 2= 1000/5= 200 units/ml

Concentration of Tube 3= 200/5= 40 units/ml

Concentration of Tube 4= 40/5= 8 units /ml

Concentration of Tube 5=8/5= 1.6 units/ ml

Concentration of Tube 6= 1.6/5= 0.32 units/ml

Concentration of Tube 7= 0.32/5= 0.064 units/ml

Minimal inhibitory concentration of a drug is the minimum concentration that will inhibit the growth of the microorganism tested after overnight incubation.

In this experiment, tube 7 showed growth of the organism. The minimum concentration of penicillin that was inhibiting growth was tube 6.

Hence, MIC of Penicillin in this assay is 0.32 units/ml.

6. a) Tube no 5 has 58 colonies. 1 ml was plated; total number of colonies per ml is 58.

This plate was a diluted 1:10 four times, as tube 1 is original undiluted culture. Hence, dilution factor is 10 X 10 X 10 X 10= 10000.

Total number of bacteria in the original culture= Number of colonies X dilution factor

                                                                                        = 58 X 10000=5.8 X 105 cells/ml.

b. There were 58 colonies in tube 4 of 1: 10 dilution series.

Dilution factor is 1000.

Total number of bacteria in original culture= 58 X 1000= 5.8 X 104 cells/ml.

c. Dilution factor in tube 5 = 10000.

0.1 ml was plated from the culture, which had 58 colonies. Hence, 1 ml would have 580 colonies.

Total number of bacteria in original culture= 580 X 1000= 5.8 X 106 cells/ml.

d. Dilution factor in tube 5 = 10000.

0.25 ml was plated from the culture, which had 58 colonies. Hence, 1 ml would have 58 X 4=232 colonies.

Total number of bacteria in original culture= 232 X 1000= 2.32 X 106 cells/ml.